# Recurrence in collective dynamics: From the hyperspace to fuzzy dynamical systems

### v1

@@ Abstract
We study for a dynamical system $f:X\longrightarrow X$ some of the principal topological recurrence-kind properties with respect to the induced maps $\overline{f}:\mathcal{K}(X)\longrightarrow\mathcal{K}(X)$, on the hyperspace of non-empty compact subsets of $X$, and $\hat{f}:\mathcal{F}(X)\longrightarrow\mathcal{F}(X)$, on the space of normal fuzzy sets consisting of the upper-semicontinuous functions $u:X\longrightarrow[0,1]$ with compact support and such that $u^{-1}(\{1\})\neq\varnothing$. In particular, we characterize the properties of topological and multiple recurrence for the extended systems $(\mathcal{K}(X),\overline{f})$ and $(\mathcal{F}(X),\hat{f})$. Special attention is given to the case where the underlying space is completely metrizable, for which we obtain some stronger point-recurrence equivalences.
@@ 1 Introduction
In addition, when the underlying space $X$ is metrizable, another type of collective dynamics has recently been considered: the dynamical system $(X,f)$ induces a system $(\mathcal{F}(X),\hat{f})$, on the space $\mathcal{F}(X)$ of normal fuzzy sets, where $\hat{f}:\mathcal{F}(X)\longrightarrow\mathcal{F}(X)$ is called the Zadeh extension of $f$. In this context: Jardón et al. studied in [19, 20, 21] the relations between $(X,f)$, $(\mathcal{K}(X),\overline{f})$ and $(\mathcal{F}(X),\hat{f})$ with respect to some expansive and transitivity properties; Cánovas and Kupka in [7], and Kim et al. in [23, 32] focused on the notion of topological entropy; and other dynamical properties, including the so-called Devaney, Li-Yorke and distributional chaos, were studied in [25, 30, 33].
@@
The paper is organized as follows. In Section 2 we introduce the notation used, together with the general background needed. In Section 3 we focus on the recurrence of the system $(\mathcal{K}(X),\overline{f})$ where $X$ is a topological space, and in Section 4 we assume that $X$ is a metric space in order to compare the recurrence of $(\mathcal{K}(X),\overline{f})$ and $(\mathcal{F}(X),\hat{f})$. Finally, in Section 5 we let $X$ be a complete metric space to obtain stronger dynamical results in terms of various point-recurrence properties, and in Section 6 we include some conclusions and we leave an open problem.
@@ 2 Notation and general background
In this paper we combine the theory of dynamical systems with that of hyperspaces of compact and fuzzy sets, and we recall in this section the basic definitions and notation that we use through the whole paper. We start by the dynamical properties that we are about to study, and then we include some basic facts about the topologies that we consider on the hyperspaces $\mathcal{K}(X)$ and $\mathcal{F}(X)$.
From now on we let $\mathbb{N}$ be the set of strictly positive integers and we will write $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$.
@@ 2.1 Dynamical properties: From transitivity to recurrence
  2. a dynamical system $(X,f)$ is called topologically $\mathcal{A}$-transitive, for a Furstenberg family $\mathcal{A}\subset\mathscr{P}(\mathbb{N}_{0})$, if given any pair of non-empty open subsets $U,V\subset X$ we have that the return set from $U$ to $V$, which will be denoted by
    $\mathcal{N}_{f}(U,V):=\{n\in\mathbb{N}_{0}\ ;\ f^{n}(U)\cap V\neq\varnothing\},$
    belongs to the family $\mathcal{A}$.
Typical explicit examples of this last definition arise when we consider: the family $\mathcal{A}_{\infty}$, formed by the infinite subsets of $\mathbb{N}_{0}$, so that topological $\mathcal{A}_{\infty}$-transitivity is exactly topological transitivity; the family of co-finite sets $\mathcal{A}_{cf}$, whose respective topological $\mathcal{A}_{cf}$-transitivity notion is called topological mixing; or the family $\mathcal{A}_{th}$, formed by the so-called thick sets (i.e. the sets $A\subset\mathbb{N}_{0}$ such that for every $\ell\in\mathbb{N}$ there is $n\in A$ with $[\![n,n+\ell]\!]\subset A$), and whose respective topological $\mathcal{A}_{th}$-transitivity notion coincides with the well-known property of topological weak-mixing (see for instance [16, Theorem 1.54]).
@@
The previous notion of recurrence are important in Topological Dynamics. Indeed:
@@
    The multiply recurrent dynamical systems have been deeply studied in the literature (see for instance [12, Chapter 2, Section 2]), and a they have recently been considered in the works [8, 11], but also in the recent 2017 paper [26] under the name of Van der Waerden systems.
@@
    As it happens for transitivity, the notion of “topological recurrence” can be expressed as that of “topological $\mathcal{A}_{\infty}$-recurrence” for the family $\mathcal{A}_{\infty}$ of infinite subsets of $\mathbb{N}_{0}$. However, there is no family $\mathcal{A}\subset\mathscr{P}(\mathbb{N}_{0})$ for which “topological $\mathcal{A}$-recurrence” equals “multiple recurrence”. Indeed, there exist dynamical systems that even are topologically $\mathcal{A}_{cf}$-transitive (i.e. topologically mixing) while they are not multiply recurrent (see [26, Proposition 4.17]).
@@ 2.2 The hyperspaces $\mathcal{K}(X)$ and $\mathcal{F}(X)$
We will denote by $\mathcal{B}_{H}(K,\varepsilon)$ the open ball centred at $K\in\mathcal{K}(X)$ and of radius $\varepsilon>0$ for the metric $d_{H}$, and along the paper we will need the following well-known fact:
@@
It follows from [20, Propositions 3.1 and 4.9] that, when $X$ is a Hausdorff topological space and the map $f:X\longrightarrow X$ is continuous, then the respective Zadeh extension $\hat{f}$ (defined as in (1) and restricted to the space of normal fuzzy sets $\mathcal{F}(X)$) is a well-defined self-map of $\mathcal{F}(X)$. In order to talk about continuity for $\hat{f}:\mathcal{F}(X)\longrightarrow\mathcal{F}(X)$ we must endow $\mathcal{F}(X)$ with some topology, and this forces us to consider metric spaces. Following [19, 23, 30, 32], given a metric space $(X,d)$ we will consider:
@@
    where $d_{H}$ denotes the already introduced Hausdorff metric on $\mathcal{K}(X)$. We denote by $\mathcal{B}_{\infty}(u,\varepsilon)$ the open ball centred at $u\in\mathcal{F}(X)$ and of radius $\varepsilon>0$ for the metric $d_{\infty}$, and we will denote by $\mathcal{F}_{\infty}(X)$ the topological space $(\mathcal{F}(X),d_{\infty})$ for short, which is a metric space that is non-separable as soon as $X$ has more than one point (see Section 5 for more on the non-separability of $\mathcal{F}_{\infty}(X)$).
@@
    where $\mathcal{T}$ is the set of strictly increasing homeomorphisms of the unit interval $\xi:[0,1]\longrightarrow[0,1]$. We denote by $\mathcal{B}_{0}(u,\varepsilon)$ the open ball centred at $u\in\mathcal{F}(X)$ and of radius $\varepsilon>0$ for the metric $d_{0}$, and we will denote by $\mathcal{F}_{0}(X)$ the topological space $(\mathcal{F}(X),d_{0})$ for short, which is a metric space.
It was showed in [20, Theorems 4.7 and 4.8] that $\hat{f}$ is continuous in $\mathcal{F}_{\infty}(X)$ and $\mathcal{F}_{0}(X)$ if and only if the map $f$ is continuous on the metric space $(X,d)$. Let us include some important well-known properties of the Zadeh extension $\hat{f}$ that we will use without citing them (see [9, 19, 20]):
@@ 2. Case 2: There exists some $n_{0}\in E$ such that for every compact set
    If the Furstenberg family $\mathcal{A}\subset\mathscr{P}(\mathbb{N}_{0})$ considered in the statement of Theorem 3.1 is not a filter, then the $N$-fold direct product hypothesis for every $N\in\mathbb{N}$ in statement (i) of Theorem 3.1 is necessary. Indeed, for the case of usual topological recurrence (i.e. topological $(1,\mathcal{A}_{\infty})$-recurrence), it was shown in [2, Example 4] and [15, Theorem 3.2] that given any $N\in\mathbb{N}$ there exist many dynamical systems $(X,f)$ such that:
@@ 4 Recurrence in $\mathcal{F}(X)$
As mentioned in Remark 3.2 but this time for Corollary 4.2, this result is not true for the notion of topological recurrence (neither for multiple recurrence) if we do not assume for the continuous map $f$ the $N$-fold direct product hypothesis for every $N\in\mathbb{N}$. Indeed, the already mentioned dynamical system $f:X\longrightarrow X$ constructed in [15, Theorem 3.2] for each $N\in\mathbb{N}$, acts on a metric space $X$ and fulfills that $f_{(N)}$ is topologically recurrent (and even multiply recurrent), but also that $f_{(N+1)}$ is not topologically recurrent (neither multiply recurrent).
@@ 5 The case of complete metric spaces
In this section we assume that $X$ is a complete metric space. This assumption is usually fulfilled by the phase spaces considered in the literature ($\mathbb{R}^{N}$ or $\mathbb{C}^{N}$, compact metric spaces, etc.), and it allows us to get results stronger than Corollary 4.2 by using Baire category arguments. Let us start by recalling which kind of point-recurrence notions we are going to consider.
@@
It is not hard to show that every point-recurrent dynamical system is topologically recurrent, and that every $\mathcal{AP}$-recurrent system is multiply recurrent. The converse implications are not true in general (see for instance [16, Example 12.9]), but hold when $X$ is a complete metric space by using some Baire category arguments (see [10, Proposition 2.1] and [26, Lemma 4.8]). Thus, we have the next:
@@
If the metric space $(X,d)$ is complete then we have: that the metric space $(\mathcal{K}(X),d_{H})$ is complete (see for instance [18, Exercise 2.15]); that $\mathcal{F}_{\infty}(X)\equiv(\mathcal{F}(X),d_{\infty})$ is complete (see [24]); and that, even though $\mathcal{F}_{0}(X)\equiv(\mathcal{F}(X),d_{0})$ is not necessarily complete (see [22, Theorem 3.4]), by the result [22, Theorem 3.8] (and its generalization to arbitrary metric spaces [17, Proposition 4.6]) there exists an equivalent metric $d_{1}$ (i.e. the topology induced by $d_{0}$ and $d_{1}$ coincide) for which $(\mathcal{F}(X),d_{1})$ is a complete metric space (see [22, Theorem 3.9]). The result is now a direct consequence of the fact that point-recurrence (resp. $\mathcal{AP}$-recurrence) implies topological recurrence (resp. multiple recurrence), together with Corollary 4.2 and the completeness of the spaces involved. ∎
@@
As in Remark 3.2 and at the end of Section 4, the conclusion of Corollary 5.1 is not true for the notion of point-recurrence (neither for $\mathcal{AP}$-recurrence) if we do not assume for the map $f$ the $N$-fold direct product hypothesis for every $N\in\mathbb{N}$. Indeed, the already mentioned dynamical systems constructed in [15, Theorem 3.2] are again a counterexample in this case. However, if apart from the completeness of $(X,d)$ we also assume separability, then we can delete the $N$-fold direct product hypothesis by using the notion of quasi-rigidity recently introduced in [15]:
@@
The notion of quasi-rigidity has been used in many concrete recurrence problems (see [15, 28, 29]), and it is not hard to check that every quasi-rigid system $(X,f)$ fulfills that $(X^{N},f_{(N)})$ is point-recurrent for every $N\in\mathbb{N}$. The converse holds when $X$ is a separable complete metric (also called Polish) space by the Baire category theorem (see [15, Theorem 2.5]), and we can obtain the next:
@@
If $(X,d)$ is separable then we have: that $(\mathcal{K}(X),d_{H})$ is separable (see for instance [18, Exercise 1.18]); and that $\mathcal{F}_{0}(X)\equiv(\mathcal{F}(X),d_{0})$ is separable (see [20, Theorem 4.12]). Since these spaces are also completely metrizable (see the proof of Corollary 5.1 for details) the equivalences (i) $\Leftrightarrow$ (ii) $\Leftrightarrow$ (iv) follow from the fact that quasi-rigidity implies topological recurrence for every $N$-fold direct product, together with Corollary 4.2, separability and completeness.
@@
6 Conclusions
We have examined the interaction between various individual recurrence-kind properties presented by a discrete dynamical system $(X,f)$ and the respective collective dynamics of its extensions $(\mathcal{K}(X),\overline{f})$ and $(\mathcal{F}(X),\hat{f})$. In particular, we have proved that the topological $(\ell,\mathcal{A})$-recurrence for every $(X^{N},f_{(N)})$ is equivalent to the fact that $(\mathcal{K}(X),\overline{f})$ has this same property when $X$ is any topological space, but also equivalent to the topological $(\ell,\mathcal{A})$-recurrence of $(\mathcal{F}(X),\hat{f})$ with respect to the metrics $d_{\infty}$ and $d_{0}$ when $(X,d)$ is a metric space. As a consequence, the same results hold: for the particular notions of topological recurrence, multiple recurrence and topological $\mathcal{A}$-recurrence in general; but also for the stronger notions of point-recurrence and point-$\mathcal{AP}$-recurrence (also quasi-rigidity) when $X$ is a complete (and also separable) metric space.
We would like to close the paper by mentioning a possible future line of research. Indeed, having in mind the works [5] and [34] about collective dynamics in the context of Linear Dynamics, one can consider the following situation:
  1. Given a continuous linear operator $T:X\longrightarrow X$ let $\mathcal{C}(X)$ be the hyperspace of non-empty compact and convex subsets of $X$. It is well-known that the hyperextension map $\overline{T}:\mathcal{K}(X)\longrightarrow\mathcal{K}(X)$ leaves the set $\mathcal{C}(X)$ invariant by the continuity and linearity of $T$. Moreover, considering $\mathcal{C}(X)$ as a topological subspace of $\mathcal{K}(X)$ with the restricted Vietoris topology, one can study the dynamics of the restriction
    $\overline{T}|_{\mathcal{C}(X)}:=\tilde{T}:\mathcal{C}(X)\longrightarrow\mathcal{C}(X),$
    where $\tilde{T}(K):=\overline{T}(K)=T(K)=\{Tx\ ;\ x\in K\}$ for every set $K\in\mathcal{C}(X)$.
The following natural open question arises (see [5] for more details):

### v2

@@ Abstract
We study for a dynamical system $f:X\longrightarrow X$ some of the principal topological recurrence-kind properties with respect to the induced maps $\overline{f}:\mathcal{K}(X)\longrightarrow\mathcal{K}(X)$, on the hyperspace of non-empty compact subsets of $X$, and $\hat{f}:\mathcal{F}(X)\longrightarrow\mathcal{F}(X)$, on the space of normal fuzzy sets consisting of the upper-semicontinuous functions $u:X\longrightarrow[0,1]$ with compact support and such that $u^{-1}(\{1\})\neq\varnothing$. In particular, we characterize the properties of topological and multiple recurrence for the extended systems $(\mathcal{K}(X),\overline{f})$ and $(\mathcal{F}(X),\hat{f})$, which cover the cases of the so-called nonwandering and Van der Waerden systems. Special attention is given to the case where the underlying space is completely metrizable, for which we obtain some stronger point-recurrence equivalences.
@@ 1 Introduction
In addition, when the underlying space $X$ is metrizable, another type of collective dynamics has recently been considered: the dynamical system $(X,f)$ induces a system $(\mathcal{F}(X),\hat{f})$, on the space $\mathcal{F}(X)$ of normal fuzzy sets, where $\hat{f}:\mathcal{F}(X)\longrightarrow\mathcal{F}(X)$ is called the Zadeh extension of $f$. In this context: Jardón et al. studied in [20, 21, 22] the relations between $(X,f)$, $(\mathcal{K}(X),\overline{f})$ and $(\mathcal{F}(X),\hat{f})$ with respect to some expansive and transitivity properties; Cánovas and Kupka in [7], and Kim et al. in [24, 34] focused on the notion of topological entropy; and other dynamical properties, including the so-called Devaney, Li-Yorke and distributional chaos, were studied in [26, 32, 35, 36].
@@
The paper is organized as follows. In Section 2 we introduce the notation together with the general background needed. We recall there both the dynamical properties considered, including nonwandering and Van der Waerden systems (see Remark 2.1 below), and the specific topological setting that we use for the hyperspaces $\mathcal{K}(X)$ and $\mathcal{F}(X)$. In Section 3 we focus on the system $(\mathcal{K}(X),\overline{f})$ where $X$ is a topological space, in Section 4 we assume that $X$ is a metric space to compare the recurrence of $(\mathcal{K}(X),\overline{f})$ with that of $(\mathcal{F}(X),\hat{f})$, and in Section 5 we let $X$ be a complete metric space to obtain stronger dynamical results in terms of various point-recurrence properties. In Section 6 we include some open problems asking whether some of the complete spaces results still hold in the non-complete case, also if our main results (Theorems 3.1 and 4.1) can be extended to linear dynamical systems, and we finally discuss the possibility of considering other topologies on $\mathcal{K}(X)$ and $\mathcal{F}(X)$ as in [35].
@@ 2 Notation and general background
In this paper we combine the theory of dynamical systems with that of hyperspaces of compact and fuzzy sets, and we recall in this section the basic definitions and notation that we use through the whole paper. We start by the dynamical properties that we are about to study, and then we include some basic facts about the topologies that we consider on the hyperspaces $\mathcal{K}(X)$ and $\mathcal{F}(X)$. From now on we let $\mathbb{N}$ be the set of strictly positive integers and we will write $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$.
@@ 2.1 Dynamical properties: From transitivity to recurrence
  2. a dynamical system $(X,f)$ is called topologically $\mathcal{A}$-transitive, for a Furstenberg family $\mathcal{A}\subset\mathscr{P}(\mathbb{N}_{0})$, if given any pair of non-empty open subsets $U,V\subset X$ we have that the return set from $U$ to $V$, which will be denoted by $\mathcal{N}_{f}(U,V):=\{n\in\mathbb{N}_{0}\ ;\ f^{n}(U)\cap V\neq\varnothing\}$, belongs to the family $\mathcal{A}$.
Typical explicit examples of this last definition arise when we consider: the family $\mathcal{A}_{\infty}$, formed by the infinite subsets of $\mathbb{N}_{0}$, so that topological $\mathcal{A}_{\infty}$-transitivity is exactly topological transitivity; the family of co-finite sets $\mathcal{A}_{cf}$, whose respective topological $\mathcal{A}_{cf}$-transitivity notion is called topological mixing; or the family $\mathcal{A}_{th}$, formed by the so-called thick sets (i.e. the sets $A\subset\mathbb{N}_{0}$ such that for every $\ell\in\mathbb{N}$ there is $n\in A$ with $[\![n,n+\ell]\!]\subset A$), and whose respective topological $\mathcal{A}_{th}$-transitivity notion coincides with the well-known property of topological weak-mixing (see for instance [17, Theorem 1.54]).
@@
The previous notions of recurrence are important in Topological Dynamics. Indeed:
@@
    The multiply recurrent dynamical systems have been deeply studied in the literature (see for instance [12, Chapter 2, Section 2]), and they have recently been considered in the works [8, 11], but also in the recent 2017 paper [27] under the name of Van der Waerden systems.
@@
    As it happens for transitivity, the notion of “topological recurrence” can be expressed as that of “topological $\mathcal{A}_{\infty}$-recurrence” for the family $\mathcal{A}_{\infty}$ of infinite subsets of $\mathbb{N}_{0}$. However, there is no family $\mathcal{A}\subset\mathscr{P}(\mathbb{N}_{0})$ for which “topological $\mathcal{A}$-recurrence” equals “multiple recurrence”. Indeed, there exist dynamical systems that even are topologically $\mathcal{A}_{cf}$-transitive (i.e. topologically mixing) while they are not multiply recurrent (see [27, Proposition 4.17]).
@@ 2.2 The hyperspaces $\mathcal{K}(X)$ and $\mathcal{F}(X)$
We will denote by $\mathcal{B}_{H}(K,\varepsilon)$ the open ball centered at $K\in\mathcal{K}(X)$ and of radius $\varepsilon>0$ for the metric $d_{H}$, and along the paper we will need the following well-known fact:
@@
It follows from [21, Propositions 3.1 and 4.9] that, when $X$ is a Hausdorff topological space and the map $f:X\longrightarrow X$ is continuous, then the respective Zadeh extension $\hat{f}$ (defined as in (1) and restricted to the space of normal fuzzy sets $\mathcal{F}(X)$) is a well-defined self-map of $\mathcal{F}(X)$. In order to talk about continuity for $\hat{f}:\mathcal{F}(X)\longrightarrow\mathcal{F}(X)$ we must endow $\mathcal{F}(X)$ with some topology, and this forces us to consider metric spaces. Following [20, 24, 32, 34], given a metric space $(X,d)$ we will consider:
@@
    where $d_{H}$ denotes the already introduced Hausdorff metric on $\mathcal{K}(X)$. We denote by $\mathcal{B}_{\infty}(u,\varepsilon)$ the open ball centered at $u\in\mathcal{F}(X)$ and of radius $\varepsilon>0$ for the metric $d_{\infty}$, and we will denote by $\mathcal{F}_{\infty}(X)$ the topological space $(\mathcal{F}(X),d_{\infty})$ for short, which is a metric space that is non-separable as soon as $X$ has more than one point (see Section 5 for more on the non-separability of $\mathcal{F}_{\infty}(X)$).
@@
    where $\mathcal{T}$ is the set of strictly increasing homeomorphisms of the unit interval $\xi:[0,1]\longrightarrow[0,1]$. We denote by $\mathcal{B}_{0}(u,\varepsilon)$ the open ball centered at $u\in\mathcal{F}(X)$ and of radius $\varepsilon>0$ for the metric $d_{0}$, and we will denote by $\mathcal{F}_{0}(X)$ the topological space $(\mathcal{F}(X),d_{0})$ for short, which is a metric space.
It was showed in [21, Theorems 4.7 and 4.8] that $\hat{f}$ is continuous in $\mathcal{F}_{\infty}(X)$ and $\mathcal{F}_{0}(X)$ if and only if the map $f$ is continuous on the metric space $(X,d)$. Let us include some important well-known properties of the Zadeh extension $\hat{f}$ that we will use without citing them (see [9, 20, 21]):
@@ 2. Case 2: There exists some $n_{0}\in E$ such that for every compact set
    If the Furstenberg family $\mathcal{A}\subset\mathscr{P}(\mathbb{N}_{0})$ considered in the statement of Theorem 3.1 is not a filter, then the $N$-fold direct product hypothesis for every $N\in\mathbb{N}$ in statement (i) of Theorem 3.1 is necessary. Indeed, for the case of usual topological recurrence (i.e. topological $(1,\mathcal{A}_{\infty})$-recurrence), it was shown in [2, Example 4] and [16, Theorem 3.2] that given any $N\in\mathbb{N}$ there exist many dynamical systems $(X,f)$ such that:
@@ 4 Recurrence in $\mathcal{F}(X)$
As mentioned in Remark 3.2 but this time for Corollary 4.2, this result is not true for the notion of topological recurrence (neither for multiple recurrence) if we do not assume for the continuous map $f$ the $N$-fold direct product hypothesis for every $N\in\mathbb{N}$. Indeed, the already mentioned dynamical system $f:X\longrightarrow X$ constructed in [16, Theorem 3.2] for each $N\in\mathbb{N}$, acts on a metric space $X$ and fulfills that $f_{(N)}$ is topologically recurrent (and even multiply recurrent), but also that $f_{(N+1)}$ is not topologically recurrent (neither multiply recurrent).
@@ 5 The case of complete metric spaces
In this section we assume that $X$ is a complete metric space. This assumption is usually fulfilled by the phase spaces considered in the literature ($\mathbb{R}^{N}$ or $\mathbb{C}^{N}$, compact metric spaces, etc.), and it allows us to get stronger results than Corollary 4.2 by using Baire category arguments. Let us start by recalling which kind of point-recurrence notions we are going to consider.
@@
It is not hard to show that every point-recurrent dynamical system is topologically recurrent, and that every $\mathcal{AP}$-recurrent system is multiply recurrent. The converse implications are not true in general (see for instance [17, Example 12.9]), but hold when $X$ is a complete metric space by using some Baire category arguments (see [10, Proposition 2.1] and [27, Lemma 4.8]). Thus, we have the next:
@@
If the metric space $(X,d)$ is complete then we have: that the metric space $(\mathcal{K}(X),d_{H})$ is complete (see for instance [19, Exercise 2.15]); that $\mathcal{F}_{\infty}(X)\equiv(\mathcal{F}(X),d_{\infty})$ is complete (see [25]); and that, even though $\mathcal{F}_{0}(X)\equiv(\mathcal{F}(X),d_{0})$ is not necessarily complete (see [23, Theorem 3.4]), by the result [23, Theorem 3.8] (and its generalization to arbitrary metric spaces [18, Proposition 4.6]) there exists an equivalent metric $d_{1}$ (i.e. the topology induced by $d_{0}$ and $d_{1}$ coincide) for which $(\mathcal{F}(X),d_{1})$ is a complete metric space (see [23, Theorem 3.9]). The result is now a direct consequence of the fact that point-recurrence (resp. $\mathcal{AP}$-recurrence) implies topological recurrence (resp. multiple recurrence), together with Corollary 4.2 and the completeness of the spaces involved. ∎
@@
As in Remark 3.2 and at the end of Section 4, the conclusion of Corollary 5.1 is not true for the notion of point-recurrence (neither for $\mathcal{AP}$-recurrence) if we do not assume for the map $f$ the $N$-fold direct product hypothesis for every $N\in\mathbb{N}$. Indeed, the already mentioned dynamical systems constructed in [16, Theorem 3.2] are again a counterexample in this case. However, if apart from the completeness of $(X,d)$ we also assume separability, then we can delete the $N$-fold direct product hypothesis by using the notion of quasi-rigidity recently introduced in [16]:
@@
The notion of quasi-rigidity has been used in many concrete recurrence problems (see [16, 30, 31]), and it is not hard to check that every quasi-rigid system $(X,f)$ fulfills that $(X^{N},f_{(N)})$ is point-recurrent for every $N\in\mathbb{N}$. The converse holds when $X$ is a separable complete metric (also called Polish) space by the Baire category theorem (see [16, Theorem 2.5]), and we can obtain the next:
@@
If $(X,d)$ is separable then we have: that $(\mathcal{K}(X),d_{H})$ is separable (see for instance [19, Exercise 1.18]); and that $\mathcal{F}_{0}(X)\equiv(\mathcal{F}(X),d_{0})$ is separable (see [21, Theorem 4.12]). Since these spaces are also completely metrizable (see the proof of Corollary 5.1 for details) the equivalences (i) $\Leftrightarrow$ (ii) $\Leftrightarrow$ (iv) follow from the fact that quasi-rigidity implies topological recurrence for every $N$-fold direct product, together with Corollary 4.2, separability and completeness.
@@
6 Conclusions and open problems
We have examined the interaction between various individual recurrence-kind properties presented by a discrete dynamical system $(X,f)$ and the respective collective dynamics of its extensions $(\mathcal{K}(X),\overline{f})$ and $(\mathcal{F}(X),\hat{f})$. In particular, we have proved that the topological $(\ell,\mathcal{A})$-recurrence for every $(X^{N},f_{(N)})$ is equivalent to the fact that $(\mathcal{K}(X),\overline{f})$ has this same property when $X$ is any topological space, but also equivalent to the topological $(\ell,\mathcal{A})$-recurrence of $(\mathcal{F}(X),\hat{f})$ with respect to the metrics $d_{\infty}$ and $d_{0}$ when $(X,d)$ is a metric space. As a consequence, the same results hold: for the particular notions of topological recurrence, multiple recurrence and topological $\mathcal{A}$-recurrence in general; but also for the stronger notions of point-recurrence and point-$\mathcal{AP}$-recurrence (also quasi-rigidity) when $X$ is a complete (and also separable) metric space. Now, we would like to close the paper by mentioning some open problems and possible future lines of research.
First of all, and following Section 5, we would like to highlight in a proper way the left implications mentioned in Remarks 5.2 and 5.4. Indeed, we have proved there that some of the implications and equivalences stated in Corollaries 5.1 and 5.3 still hold when the underlying space is not complete. Answering the next open question would extend these results to a much more general class of spaces: